# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

### v2

@@ H.1 Lower Bound for Q-value Learning
and
$\alpha=8\varepsilon\frac{|\beta|}{\gamma}\frac{1}{e^{|\beta|\frac{1}{1-\gamma}}-1},$
for any $\varepsilon$ in the range $\varepsilon<\frac{1}{40}\frac{\gamma}{|\beta|}(1-e^{-|\beta|})$ .
@@
In particular, this means that the events $B_{0}:=\{|Q^{*}_{M_{0}}(z_{i})-Q_{t}^{\mathcal{U}}(z_{i})|\leq\varepsilon\}$ and $B_{1}:=\{|Q^{*}_{M_{1}}(z_{i})-Q_{t}^{\mathcal{U}}(z_{i})|\leq\varepsilon\}$ are disjoint events Let $t$ be the number of times the algorithm tries $z_{i}$. Since $\mathcal{U}$ is $(\varepsilon,\delta)$-correct it holds that $\mathbb{P}_{0}(B_{0})\geq 1-\delta\geq\frac{3}{4}$.
Let $k$ be the number of transitions from $z_{i}$ to $z_{i}^{G}$ in the t trials. We then define $\tilde{k},\tilde{p}$ and $\theta$ by
$\tilde{k}:=\begin{cases}k\quad&\text{if }p\geq\frac{1}{2}\\ t-k\quad&\text{if }p<\frac{1}{2}\end{cases}$
$\theta:=\exp\big(-\frac{32\alpha^{2}t}{p(1-p)}\big)$
$\tilde{p}=\max\{p,1-p\}$
@@
$\mathcal{E}$ $=\bigg\{\tilde{p}t-\tilde{k}\leq\sqrt{2p(1-p)t\log(\frac{8}{2\theta})}\bigg\}$
for which we have $\mathbb{P}_{0}(\mathcal{E})>\frac{3}{4}$ by Lemma 16 in [23] and thus $\mathbb{P}_{0}(B_{0}\cap\mathcal{E})>\frac{1}{2}$. Now by Theorem 9 we get that
$\mathbb{P}_{1}(B_{0})\geq\mathbb{P}_{1}(B_{0}\cap\mathcal{E})=\mathbb{E}_{1}[1_{\mathcal{E}}1_{B_{0}}]=\mathbb{E}_{0}\bigg[\frac{L_{1}}{L_{0}}1_{\mathcal{E}}1_{B_{0}}\bigg]\geq\frac{\theta}{4}\mathbb{E}_{0}[1_{\mathcal{E}}1_{B_{0}}]=\frac{\theta}{4}\mathbb{P}_{0}(\mathcal{E}\cap B_{0})\geq\frac{\theta}{8}\,.$
@@
Next we use the fact that the structure of the MDPs is such that information on the $Q$-value of any state-action pair in $Z$ carries no information on the $Q$-values of any other state-action pair in $Z$.
Let $n:=SA$. If the number of total transition samples is less than $\frac{n}{2}\tilde{T}(\varepsilon,\delta)$ there must be at least $n/2$ state-action pairs $z_{i}$ that has been tried no more than $\tilde{T}(\varepsilon,\delta)$ times which without loss of generality we might assume are the state-action pairs $\{z_{i}\}_{i=1}^{n/2}$.
@@
$\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}\cap\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq n/2})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{n/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}_{1}(\{T_{i}=t_{i}\}_{1\leq i\leq n/2})\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}\cap\{T_{i}=t_{i}\}_{1\leq i\leq n/2})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{n/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}_{1}(\{T_{i}=t_{i}\}_{1\leq i\leq n/2})\prod_{1\leq i\leq n/2}\mathbb{P}_{1}(\Lambda_{i}^{c}\cap\{T_{i}=t_{i}\})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{n/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}_{1}(\{T_{i}=t_{i}\}_{1\leq i\leq n/2})(1-\delta)^{n/2},$
@@
$\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}|\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq n/2})\leq(1-\delta)^{\frac{n}{2}}\,.$
Thus, if the total number of transitions $T$ is less than $\frac{n}{2}\tilde{T}(\varepsilon,\delta)$, then
@@
$\geq 1-\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}|\{T_{z_{i}}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq n/2})$
@@
$T=\frac{(S^{\prime}-2)A}{4096}\frac{\gamma^{2}}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\log(\frac{(S^{\prime}-2)A}{32\delta})$ (71)
@@
H.2 Lower Bound for Policy Learning
For a lower bound we construct the following class of MDPs with $S^{\prime}:=S+2$ states and $A^{\prime}:=A+1$ actions where the first states are labelled $s_{1},\ldots,s_{S},s^{G},s^{B}$ and the actions are labelled $a_{0},a_{1},...,a_{A}$. The states $s^{G}$ and $s^{B}$ are absorbing under any actions and $R(s^{G},a)=1$ for all $j$ and $R(s^{B},a)=0$ for all $a\in A$. For the states $s\in\{s_{1},...,s_{S}\}$, we have that $R(s,a)=0$ for all $a\in A$.
@@ H.2 Lower Bound for Policy Learning
where we allow for
$0<\varepsilon<\frac{\gamma}{50|\beta|}\big(1-e^{-|\beta|\frac{1}{1-\gamma}}\big)\,$,
which ensures that $\alpha\leq\frac{e^{-|\beta|\frac{1}{1-\gamma}}}{10}$.
@@
$\frac{5}{21}\frac{\gamma}{|\beta|}\alpha(e^{|\beta|\frac{1}{1-\gamma}}-1)$
@@
Now haven shown that all non-optimal actions are $\varepsilon$-bad, we wish to show that any algorithm that is $(\varepsilon,\delta)$-correct on $H_{0}^{i}$, i.e. choosing the action $a_{0}$ with probability at least $1-\delta$, will also have a probability of choosing $a_{0}$ on $H_{l}^{i}$ that is larger than $\delta$ provided that $a_{l}$ is not tried sufficiently many times under $H_{0}^{i}$.
@@
Form the Chernoff-Hoeffding bound and as shown in [23] we have that $\mathbb{P}_{0}(\mathcal{E)}>\frac{3}{4}$ and so $\mathbb{P}_{0}(B\cap\mathcal{E)}>\frac{1}{2}$. From Theorem 9, we get that
$\mathbb{P}_{1}(B)\geq\mathbb{P}_{1}(B\cap\mathcal{E)}=\mathbb{E}_{1}[1_{B}1_{\mathcal{E}}]\geq\mathbb{E}_{0}\bigg[\frac{L_{1}(W)}{L_{0}(W)}1_{\mathcal{E}}1_{B}\bigg]\geq\mathbb{E}_{0}\bigg[\frac{\theta}{4}1_{\mathcal{E}}1_{B}\bigg]=\frac{\theta}{4}\mathbb{P}_{0}(\mathcal{E}\cap B)\geq\frac{\theta}{8}\,.$ (73)
@@
$t<\tilde{t}(\varepsilon,\delta):=\frac{1}{800}\log(\frac{1}{8\delta})\frac{\gamma^{2}}{\varepsilon^{2}}\cdot\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}$ (74)
@@
Since this holds for all the $A$ hypotheses $H_{l}^{i},l=1,2,...,A$ it follows that the algorithm needs at least $\tilde{T}(\varepsilon,\delta):=A\tilde{t}(\varepsilon,\delta)$ samples to be $(\varepsilon,\delta)$-correct on the state $s_{i}$.
Next we use the fact that the structure of the MDPs is such that information used to determine $\pi^{*}(s_{i})$ carries no information to determine $\pi^{*}(s_{j})$ for $i\neq j$.
[…]
@@
Thus, by defining the events $\Lambda_{i}:=\{|V_{M_{1}}^{*}(s_{i})-V^{\pi^{\mathcal{U}}_{T_{i}}}(s_{i})|>\varepsilon\}$ we have that $\Lambda_{i}$ and $\Lambda_{j}$ are conditionally independent given $T_{i}$ and $T_{j}$. We then have that for the MDP $M_{1}\in\mathbb{M}$ (The one corresponding to the hypothesis $H_{1}:=\{H_{1}^{i}|1\leq i\leq n\}$) it holds that
$\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}\cap\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/2})\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}\cap\{T_{i}=t_{i}\}_{1\leq i\leq S/2})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/2})\prod_{1\leq i\leq S/2}\mathbb{P}(\Lambda_{i}^{c}\cap\{T_{i}=t_{i}\})$
$=\sum_{t_{1}=0}^{\tilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\tilde{T}(\varepsilon,\delta)}\mathbb{P}(\{T_{i}=t_{i}\}_{1\leq i\leq S/2})(1-\delta)^{S/2}\,$,
$\mathbb{P}\big(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}|\{T_{i}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2}\big)\leq(1-\delta)^{\frac{S}{2}}\,.$
Thus, if the total number of transitions $T$ is less than $\frac{S}{2}\tilde{T}(\varepsilon,\delta)$ on the MDP $M_{0}$ corresponding to the hypothesis $H_{0}:\{H_{0}^{i}|1\leq i\leq n\}$, then on $M_{1}$ it holds that
@@
$\geq 1-\mathbb{P}(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}|\{T_{z_{i}}\leq\tilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2})$
@@
when $\frac{\delta S}{2}\leq 1$ by Lemma 6. By setting $\delta^{\prime}=\delta\frac{S}{4}$ and substituting back $S^{\prime}$ and $A^{\prime}$ we obtain the result. This shows that if the number of samples is smaller than
@@
on $M_{0}$ then on $M_{1}$ it holds that $\mathbb{P}(\|V^{*}-V^{\pi^{\mathcal{U}}_{T}}\|>\varepsilon)>\delta$.

### v3

@@ F.2 Lower Bound for Q-value Learning
and $\alpha=8\varepsilon\frac{|\beta|}{\gamma}\frac{1}{e^{|\beta|\frac{1}{1-\gamma}}-1}$, for any $\varepsilon$ in the range $\varepsilon<\frac{1}{40}\frac{\gamma}{|\beta|}(1-e^{-|\beta|\frac{1}{1-\gamma}})$ .
@@
In particular, this means that the events $B_{0}:=\{|Q^{*}_{M_{0}}(z_{i})-Q_{t}^{\mathcal{U}}(z_{i})|\leq\varepsilon\}$ and $B_{1}:=\{|Q^{*}_{M_{1}}(z_{i})-Q_{t}^{\mathcal{U}}(z_{i})|\leq\varepsilon\}$ are disjoint events. Let $t$ be the number of times the algorithm tries $z_{i}$. Since $\mathcal{U}$ is $(\varepsilon,\delta)$-correct it holds that $\mathbb{P}_{0}(B_{0})\geq 1-\delta\geq\frac{3}{4}$.
Let $k$ be the number of transitions from $z_{i}$ to $s^{G}$ in the t trials. We then define $\tilde{k},\tilde{p}$ and $\theta$ by
$\theta:=\exp\Big(-\frac{32\alpha^{2}t}{p(1-p)}\Big),\qquad\tilde{p}=\max\{p,1-p\},\qquad\tilde{k}:=\begin{cases}k\quad&\text{if }p\geq\frac{1}{2}\\ t-k\quad&\text{if }p<\frac{1}{2}\end{cases}$
@@
$\mathcal{E}$ $=\bigg\{\tilde{p}t-\tilde{k}\leq\sqrt{2p(1-p)t\log(\frac{8}{2\theta})}\bigg\},$
for which, we have $\mathbb{P}_{0}(\mathcal{E})>\frac{3}{4}$ by Lemma 16 in [30] and thus $\mathbb{P}_{0}(B_{0}\cap\mathcal{E})>\frac{1}{2}$. Now by Theorem 6, we get that
$\mathbb{P}_{1}(B_{0})\geq\mathbb{P}_{1}(B_{0}\cap\mathcal{E})=\mathbb{E}_{1}[\mathbbmss{1}_{\mathcal{E}}\mathbbmss{1}_{B_{0}}]=\mathbb{E}_{0}\bigg[\frac{L_{1}}{L_{0}}\mathbbmss{1}_{\mathcal{E}}\mathbbmss{1}_{B_{0}}\bigg]\geq\frac{\theta}{4}\mathbb{E}_{0}[\mathbbmss{1}_{\mathcal{E}}\mathbbmss{1}_{B_{0}}]=\frac{\theta}{4}\mathbb{P}_{0}(\mathcal{E}\cap B_{0})\geq\frac{\theta}{8}\,.$
@@
Let $n:=SA$. If the number of total transition samples is less than $\frac{n}{2}\widetilde{T}(\varepsilon,\delta)$ there must be at least $n/2$ state-action pairs $z_{i}$ that has been tried no more than $\widetilde{T}(\varepsilon,\delta)$ times which without loss of generality we might assume are the state-action pairs $\{z_{i}\}_{i=1}^{n/2}$.
@@
$\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}$ $\cap\{T_{i}\leq\widetilde{T}(\varepsilon,\delta)\}_{1\leq i\leq n/2})$
$=\sum_{t_{1}=0}^{\widetilde{T}(\varepsilon,\delta)}\dots\sum_{t_{n/2}=0}^{\widetilde{T}(\varepsilon,\delta)}\mathbb{P}_{1}(\{T_{i}=t_{i}\}_{1\leq i\leq n/2})\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}\cap\{T_{i}=t_{i}\}_{1\leq i\leq n/2})$
$=\sum_{t_{1}=0}^{\widetilde{T}(\varepsilon,\delta)}\dots\sum_{t_{n/2}=0}^{\widetilde{T}(\varepsilon,\delta)}\mathbb{P}_{1}(\{T_{i}=t_{i}\}_{1\leq i\leq n/2})\prod_{1\leq i\leq n/2}\mathbb{P}_{1}(\Lambda_{i}^{c}\cap\{T_{i}=t_{i}\})$
$=\sum_{t_{1}=0}^{\widetilde{T}(\varepsilon,\delta)}\dots\sum_{t_{n/2}=0}^{\widetilde{T}(\varepsilon,\delta)}\mathbb{P}_{1}(\{T_{i}=t_{i}\}_{1\leq i\leq n/2})(1-\delta)^{n/2},$
@@
$\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}|\{T_{i}\leq\widetilde{T}(\varepsilon,\delta)\}_{1\leq i\leq n/2})\leq(1-\delta)^{\frac{n}{2}}\,.$
Thus, if the total number of transitions $T$ is less than $\frac{n}{2}\widetilde{T}(\varepsilon,\delta)$, then
@@
$\geq 1-\mathbb{P}_{1}(\{\Lambda_{i}^{c}\}_{1\leq i\leq n/2}|\{T_{z_{i}}\leq\widetilde{T}(\varepsilon,\delta)\}_{1\leq i\leq n/2})$
@@
$T=\frac{(S^{\prime}-2)A}{4096}\frac{\gamma^{2}}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\log\Big(\frac{(S^{\prime}-2)A}{32\delta}\Big)$ (14)
@@
F.3 Lower Bound for Policy Learning
For a lower bound we construct the following class of MDPs with $S^{\prime}:=S+2$ states and $A^{\prime}:=A+1$ actions where the first states are labelled $s_{1},\ldots,s_{S},s^{G},s^{B}$ and the actions are labelled $a_{0},a_{1},...,a_{A}$. The states $s^{G}$ and $s^{B}$ are absorbing under any actions and $R(s^{G},a)=1$ for all $j$ and $R(s^{B},a)=0$ for all $a\in\mathcal{A}$. For the states $s\in\{s_{1},...,s_{S}\}$, we have that $R(s,a)=0$ for all $a\in\mathcal{A}$.
@@ F.3 Lower Bound for Policy Learning
where we allow for $0<\varepsilon<\frac{\gamma}{50|\beta|}\big(1-e^{-|\beta|\frac{1}{1-\gamma}}\big)$, which ensures that $\alpha\leq\frac{e^{-|\beta|\frac{1}{1-\gamma}}}{10}$.
@@
$=\frac{5}{21}\frac{\gamma}{|\beta|}\alpha(e^{|\beta|\frac{1}{1-\gamma}}-1)$
@@
Now having shown that all non-optimal actions are $\varepsilon$-bad, we wish to show that any algorithm that is $(\varepsilon,\delta)$-correct on $H_{0}^{i}$, i.e. choosing the action $a_{0}$ with probability at least $1-\delta$, will also have a probability of choosing $a_{0}$ on $H_{l}^{i}$ that is larger than $\delta$ provided that $a_{l}$ is not tried sufficiently many times under $H_{0}^{i}$.
@@
From the Chernoff-Hoeffding bound and as shown in [30], we have that $\mathbb{P}_{0}(\mathcal{E)}>\frac{3}{4}$, and thus, $\mathbb{P}_{0}(B\cap\mathcal{E)}>\frac{1}{2}$. From Theorem 6, we get that
$\mathbb{P}_{1}(B)\geq\mathbb{P}_{1}(B\cap\mathcal{E)}=\mathbb{E}_{1}[\mathbbmss{1}_{B}\mathbbmss{1}_{\mathcal{E}}]\geq\mathbb{E}_{0}\bigg[\frac{L_{1}(W)}{L_{0}(W)}\mathbbmss{1}_{\mathcal{E}}\mathbbmss{1}_{B}\bigg]\geq\mathbb{E}_{0}\bigg[\frac{\theta}{4}\mathbbmss{1}_{\mathcal{E}}\mathbbmss{1}_{B}\bigg]=\frac{\theta}{4}\mathbb{P}_{0}(\mathcal{E}\cap B)\geq\frac{\theta}{8}\,.$ (16)
@@
$t<\widetilde{T}(\varepsilon,\delta):=\frac{1}{800}\log(\frac{1}{8\delta})\frac{\gamma^{2}}{\varepsilon^{2}}\cdot\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}$ (17)
@@
Since this holds for all the $A$ hypotheses $H_{l}^{i},l=1,2,\ldots,A$, it follows that the algorithm needs at least $\widetilde{T}(\varepsilon,\delta):=A\widetilde{T}(\varepsilon,\delta)$ samples to be $(\varepsilon,\delta)$-correct on the state $s_{i}$.
Next we use the fact that the structure of the MDPs is such that the information used to determine $\pi^{*}(s_{i})$ carries no information to determine $\pi^{*}(s_{j})$ for $i\neq j$.
If the number of total transition samples is less than $\frac{S}{2}\widetilde{T}(\varepsilon,\delta)$, then there must be at least $\frac{S}{2}$ states in the set $\{s_{i}\}_{i=1}^{S}$ for which some action (apart from $a_{0}$) has been tried no more than $\widetilde{T}(\varepsilon,\delta)$ times. Without loss of generality, we might assume that these are the states $\{s_{i}\}_{i=1}^{S/2}$ and that it is action $a_{1}$ that has been tried out at most $\widetilde{T}(\varepsilon,\delta)$ times in each of these states.
@@
Let us define the events $\Lambda_{i}:=\{|V_{M_{1}}^{*}(s_{i})-V^{\pi^{\mathcal{U}}_{T_{i}}}(s_{i})|>\varepsilon\}$ for $i=1,\ldots,S$. Then, we have that $\Lambda_{i}$ and $\Lambda_{j}$ are conditionally independent given $T_{i}$ and $T_{j}$. We then have that for the MDP $M_{1}\in\mathbb{M}$ –the one corresponding to the hypothesis $H_{1}:=\{H_{1}^{i}|1\leq i\leq n\}$– it holds that
$\mathbb{P}\big(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}\cap\{T_{i}\leq\widetilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2}\big)$ $=\sum_{t_{1}=0}^{\widetilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\widetilde{T}(\varepsilon,\delta)}\mathbb{P}\big(\{T_{i}=t_{i}\}_{1\leq i\leq S/2}\big)\mathbb{P}\big(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}\cap\{T_{i}=t_{i}\}_{1\leq i\leq S/2}\big)$
$=\sum_{t_{1}=0}^{\widetilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\widetilde{T}(\varepsilon,\delta)}\mathbb{P}\big(\{T_{i}=t_{i}\}_{1\leq i\leq S/2}\big)\prod_{1\leq i\leq S/2}\mathbb{P}\big(\Lambda_{i}^{c}\cap\{T_{i}=t_{i}\}\big)$
$=\sum_{t_{1}=0}^{\widetilde{T}(\varepsilon,\delta)}\dots\sum_{t_{S/2}=0}^{\widetilde{T}(\varepsilon,\delta)}\mathbb{P}\big(\{T_{i}=t_{i}\}_{1\leq i\leq S/2}\big)(1-\delta)^{S/2}\,$,
where the first line follows from the law of total probability, and the second line from independence. We now have directly that
$\mathbb{P}\Big(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}\Big|\{T_{i}\leq\widetilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2}\Big)\leq(1-\delta)^{\frac{S}{2}}\,.$
Thus, if the total number of transitions $T$ is less than $\frac{S}{2}\widetilde{T}(\varepsilon,\delta)$ on the MDP $M_{0}$ corresponding to the hypothesis $H_{0}:\{H_{0}^{i}|1\leq i\leq n\}$, then on $M_{1}$ it holds that
@@
$\geq 1-\mathbb{P}\Big(\{\Lambda_{i}^{c}\}_{1\leq i\leq S/2}\,\Big|\,\{T_{z_{i}}\leq\widetilde{T}(\varepsilon,\delta)\}_{1\leq i\leq S/2}\Big)$
@@
when $\frac{\delta S}{2}\leq 1$ by Lemma 6. By setting $\delta^{\prime}=\delta\frac{S}{4}$ and substituting back $S^{\prime}$ and $A^{\prime}$, we obtain the result. This shows that if the number of samples is smaller than
@@
on $M_{0}$, then on $M_{1}$ it holds that $\mathbb{P}(\|V^{*}-V^{\pi^{\mathcal{U}}_{T}}\|>\varepsilon)>\delta$.